# Near-Optimal Sample Complexities of Divergence-based S-rectangular Distributionally Robust Reinforcement Learning

### v1

@@ 1 Introduction
While suitable for many applications, the S-rectangular formulation in DR-RL is more challenging than its SA-rectangular counterpart, both statistically and computationally, due to the possibility of randomized optimal policies. Computationally, this requires solving a full min-max problem rather than a simpler maximization. Fortunately, Ho et al. [7] proposed an efficient method for performing Bellman updates in this setting. Statistically, the challenge arises from the fact that the space of randomized policies is exponentially larger than the space of deterministic policies typically sufficient under SA-rectangularity.
@@ 2 Literature Review
S-rectangular DR-RL: To extend the flexibility of robust MDP models, S-rectangularity was introduced in Xu and Mannor [27], Wiesemann et al. [26] as an overarching theoretical framework to constrain the adversary while retaining a dynamic programming equation. Ho et al. [7] developed an efficient optimization algorithm to solve the Bellman update under this structure. On the statistical side, Yang et al. [30] provided the first sample complexity result for S-rectangular DR-RL, achieving a rate of $\widetilde{O}(|\mathcal{S}|^{2}|\mathcal{A}|^{2}(1-\gamma)^{-4}\varepsilon^{-2})$, which is suboptimal in its dependence on the number of states and actions. More recently, Clavier et al. [2] established near-optimal rates for the S-rectangular setting under general $L_{p}$ norm uncertainty sets. However, their analysis does not directly extend to divergence-based uncertainty sets.
@@ 3.1 Classical Markov Decision Processes
$v(s)=\max_{a\in A}\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})(R(s,a,s^{\prime})+\gamma v(s^{\prime})).$ (3.1)
Let $v^{*}$ be the unique solution, then any deterministic policy $\pi^{*}:\mathcal{S}\rightarrow\mathcal{A}$ with $\pi^{*}(s)\in\arg\max_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})\\ (R(s,a,s^{\prime})+\gamma v^{*}(s^{\prime}))$ will achieve the optimal value $V_{P}^{*}(s)$.
@@ 3.2 Robust MDPs and S-Rectangularity
$\mathcal{P}_{s}(f,\rho)=\left\{P_{s,a}\in\Delta(\mathcal{S})\Bigg{|}P_{s,a}\ll\overline{P}_{s,a},\sum_{s^{\prime}\in\mathcal{S},a\in\mathcal{A}}\overline{P}_{s,a}(s^{\prime})f\left(\frac{P_{s,a}(s^{\prime})}{\overline{P}_{s,a}(s^{\prime})}\right)\leq|\mathcal{A}|\rho\right\}.$ (3.2)
@@
$v(s)=\sup_{\phi\in\Delta(\mathcal{A})}\inf_{P_{s}\in\mathcal{P}_{s}}\sum_{a\in\mathcal{A}}\phi(a)\left[\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})\left(R(s,a,s^{\prime})+\gamma v(s^{\prime})\right)\right].$ (3.4)
@@
We note that value function in (3.3) assumes an adversary that fixes a controlled transition kernel over the entire control horizon, a setting commonly referred to as a static or time-homogeneous adversarial model [8, 26, 24]. This framework can be extended to more general Markovian or history-dependent adversarial models, while still preserving Markov optimality [24].
@@
$\mathcal{T}^{\pi}(v)(s):=\inf_{P\in\mathcal{P}}\left(\sum_{a\in\mathcal{A}}\pi(a|s)\left[\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})\left(R(s,a,s^{\prime})+\gamma v(s^{\prime})\right)\right]\right)$ (3.5)
for all $s\in\mathcal{S}$. The optimal DR Bellman operator is $\mathcal{T}^{*}(v)(s):=\sup_{\pi\in\Pi}\mathcal{T}^{\pi}(v)(s),\forall s\in\mathcal{S}$.
@@ 3.3 Generative Model and the Empirical Bellman Estimator
Similarly, the empirical DR Bellman operator $\hat{\mathbf{T}}^{\pi}$ is defined as in (3.5) with $\mathcal{P}$ replaced by $\mathcal{P}_{n}$. The corresponding optimal empirical DR Bellman operator is $\hat{\mathbf{T}}^{*}(v)(s):=\sup_{\pi\in\Pi}\hat{\mathbf{T}}^{\pi}(v)(s),\forall s\in\mathcal{S}$.
Equipped with these definitions, we present our strategy to estimate the optimal value of the S-rectangular robust MDP via the empirical value function. This is motivated by the fact that $V^{*}_{\mathcal{P}}=v^{*}$ where $v^{*}$ solves (3.4).
Definition 4 (Empirical Bellman Estimator).
Given divergence function $f$ and radius parameter $\rho$, let $\mathcal{P}=\mathcal{P}(f,\rho)$ and $\mathcal{P}_{n}=\mathcal{P}_{n}(f,\rho)$. We define the empirical Bellman estimator $\hat{v}$ to $V^{*}_{\mathcal{P}}$ as the unique solution to the fixed point equation $\hat{\mathbf{T}}^{*}(\hat{v})=\hat{v}$.
The rest of this paper is dedicated to theoretical analyses and numerical validation of the statistical efficiency of estimating $V^{*}_{\mathcal{P}}=v^{*}$ using $\hat{v}$. We conclude this section by introducing the following important proposition that provides an upper bound on the $l_{\infty}$ estimation error.
@@
Let $v^{*}$,$\hat{v}$ be the solution of $\mathcal{T}^{*}(v)=v$ and $\hat{\mathbf{T}}^{*}(v)=v$, respectively. Then, the estimation error is upper bounded by
@@
with probability 1.
@@ 4 Sample Complexity Bounds for the Empirical Bellman Estimator
In this section, we establish sample complexity upper bounds to achieve an absolute $\epsilon$ error in $l_{\infty}$ distance when estimating $V^{*}_{\mathcal{P}}$ using $\hat{v}$. We focus on two specific $f$-divergence uncertainty models. When $f_{\mathrm{KL}}(t)=f(t)=t\log t$, the corresponding uncertainty set $\mathcal{P}_{s}(f_{\mathrm{KL}},\rho)$ is based on the Kullback–Leibler (KL) divergence, which is widely used in the machine learning literature. Alternatively, when $f=f_{k}$ as defined in Definition 6, the resulting $f_{k}$-divergence model captures another well-studied class of uncertainty sets [4].
@@ 4.1 The Kullback-Leibler Divergence Uncertainty Set
$\mathcal{T}^{\pi}(v)(s)=\sup_{\lambda\geq 0}\left(-\lambda|\mathcal{A}|\rho-\sum_{a\in\mathcal{A}}\lambda\log\mathbb{E}_{\overline{P}_{s,a}}\left[\exp\left(-\frac{\pi(a|s)(R(s,a,S)+\gamma v(S))}{\lambda}\right)\right]\right).$ (4.1)
The KL empirical DR Bellman operator $\hat{\mathbf{T}}^{\pi}$ satisfies (4.1) with $\overline{P}_{s,a}$ replaced by $\overline{P}_{s,a,n}$.
@@
$\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\|_{\infty}$ $\leq\frac{9\|R+\gamma v\|_{\infty}}{\sqrt{n\mathfrak{p}_{\wedge}}}\sqrt{\log\left(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta\right)}.$
@@
Under the KL-divergence uncertainty set with any $\rho\in(0,\infty)$ and $n\geq 12\mathfrak{p}_{\wedge}^{-1}\log(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)$, with probability at least $1-\eta$,
@@
Therefore, under the KL-divergence, to achieve an $\epsilon$ absolute error of estimating $V_{\mathcal{P}}^{*}=v^{*}$ with $\hat{v}$ in $l_{\infty}$ norm w.h.p., we need a total of $\widetilde{O}(|S||A|(1-\gamma)^{-4}\mathfrak{p}_{\wedge}^{-1}\epsilon^{-2})$ samples from the simulator.
@@ 4.2 $f_{k}$-Divergence Uncertainty Set
$\mathcal{T}^{\pi}(v)(s)=-\sup_{\eta\in\mathbb{R}^{|\mathcal{A}|}}\left[c\left(\sum_{a\in\mathcal{A}}\mathbb{E}_{\overline{P}_{s,a}}\left[(\eta_{a}-\pi(a|s)[R(s,a,S)+\gamma v(S)])_{+}^{k^{*}}\right]\right)^{\frac{1}{k^{*}}}+\sum_{a\in\mathcal{A}}\eta_{a}\right]$
where $c=c(k,\rho,|\mathcal{A}|)=|\mathcal{A}|^{1/k}\left(k(k-1)\rho+1\right)^{1/k}$ and $(\cdot)_{+}=\max(\cdot,0)$. The $f_{k}$ empirical DR Bellman operator $\hat{\mathbf{T}}^{\pi}$ satisfies a similar equality with $\overline{P}_{s,a}$ replaced by $\overline{P}_{s,a,n}$. $\overline{P}_{s,a,n}$.
@@
Under the $f_{k}$-divergence uncertainty set with any $\rho\in(0,\infty)$, for any $v:\mathcal{S}\rightarrow\mathbb{R}$ and $n\geq 12\mathfrak{p}_{\wedge}^{-1}\log(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)$, with probability at least $1-\eta$,
$\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\|_{\infty}$ $\leq\frac{3\cdot 2^{k^{*}}k^{*}\|R+\gamma v\|_{\infty}}{\sqrt{n\mathfrak{p}_{\wedge}}}\sqrt{\log\left(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta\right)}.$
The proof of Proposition 3 is provided in Appendix D.1. This, combined with Proposition 1, implies the following error convergence bound, whose proof is deferred to Appendix D.4.
@@
Under the $f_{k}$-divergence uncertainty set with any $\rho\in(0,\infty)$ and $n\geq 12\mathfrak{p}_{\wedge}^{-1}\log(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)$, with probability at least $1-\eta$,
@@
Therefore, under the $f_{k}$-divergence for a fixed $k$, to achieve an $\epsilon$ absolute error of estimating $V_{\mathcal{P}}^{*}=v^{*}$ with $\hat{v}$ in $l_{\infty}$ norm w.h.p., we also need a total of $\widetilde{O}(|S||A|(1-\gamma)^{-4}\mathfrak{p}_{\wedge}^{-1}\epsilon^{-2})$ samples from the simulator.

### v2

@@ 1 Introduction
While suitable for many applications, the S-rectangular formulation in DR-RL is more challenging than its SA-rectangular counterpart, both statistically and computationally, due to the possibility of randomized optimal policies. Computationally, this requires solving a full min-max problem rather than a simpler maximization. Fortunately, Ho et al. [7, 8] proposed an efficient method for performing Bellman updates in this setting. Statistically, the challenge arises from the fact that the space of randomized policies is exponentially larger than the space of deterministic policies typically sufficient under SA-rectangularity.
@@ 2 Literature Review
S-rectangular DR-RL: To extend the flexibility of robust MDP models, S-rectangularity was introduced in Xu and Mannor [29], Wiesemann et al. [28] as an overarching theoretical framework to constrain the adversary while retaining a dynamic programming equation. Ho et al. [7, 8] developed an efficient optimization algorithm to solve the Bellman update within this framework. Subsequently, Kumar et al. [10] improved upon their work by proposing a faster algorithm for the $L_{p}$ uncertainty sets. On the statistical side, Yang et al. [31] provided the first sample complexity result for S-rectangular DR-RL, achieving a rate of $\widetilde{O}(|\mathcal{S}|^{2}|\mathcal{A}|^{2}(1-\gamma)^{-4}\varepsilon^{-2})$, which is suboptimal in its dependence on the number of states and actions. More recently, Clavier et al. [2] established near-optimal rates for the S-rectangular setting under general $L_{p}$ norm uncertainty sets. However, their analysis does not directly extend to divergence-based uncertainty sets.
@@ 3.1 Classical Markov Decision Processes
$v(s)=\max_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})(R(s,a,s^{\prime})+\gamma v(s^{\prime})).$ (3.1)
Let $v^{*}$ be the unique solution, then any deterministic policy $\pi^{*}:\mathcal{S}\rightarrow\mathcal{A}$ with
$\pi^{*}(s)\in\arg\max_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})(R(s,a,s^{\prime})+\gamma v^{*}(s^{\prime}))$
will achieve the optimal value $V_{P}^{*}(s)$.
@@ 3.2 Robust MDPs and S-Rectangularity
$\mathcal{P}_{s}(f,\rho)$ $=\Bigg\{P_{s,a}\in\Delta(\mathcal{S})\Bigg|P_{s,a}\ll\overline{P}_{s,a},\sum_{s^{\prime}\in\mathcal{S},a\in\mathcal{A}}\overline{P}_{s,a}(s^{\prime})f\left(\frac{P_{s,a}(s^{\prime})}{\overline{P}_{s,a}(s^{\prime})}\right)\leq|\mathcal{A}|\rho\Bigg\}.$ (3.2)
@@
$v(s)$ $=\sup_{\phi\in\Delta(\mathcal{A})}\inf_{P_{s}\in\mathcal{P}_{s}}\sum_{a\in\mathcal{A}}\phi(a)\bigg[\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})\left(R(s,a,s^{\prime})+\gamma v(s^{\prime})\right)\bigg].$ (3.4)
@@
We note that the value function in (3.3) assumes an adversary that fixes a controlled transition kernel over the entire control horizon, a setting commonly referred to as a static or time-homogeneous adversarial model [9, 28, 26]. This framework can be extended to more general Markovian or history-dependent adversarial models, while still preserving Markov optimality [26].
@@
$\mathcal{T}^{\pi}(v)(s)=\inf_{P\in\mathcal{P}}\Bigg(\sum_{a\in\mathcal{A}}\pi(a|s)\Bigg[\sum_{s^{\prime}\in\mathcal{S}}P_{s,a}(s^{\prime})\left(R(s,a,s^{\prime})+\gamma v(s^{\prime})\right)\Bigg]\Bigg)$ (3.5)
for all $s\in\mathcal{S}$. The optimal DR Bellman operator is $\mathcal{T}^{*}(v)(s):=\max_{\pi\in\Pi}\mathcal{T}^{\pi}(v)(s)$ where the $\max$ is achieved by the same $\pi^{*}$ for all $s\in S$.
@@ 3.3 Generative Model and the Empirical Bellman Estimator
The empirical value function $V_{\hat{\mathcal{P}}}^{\pi}$ is defined in (3.3) with $\mathcal{P}$ replaced by $\mathcal{P}_{n}$. The empirical DR Bellman operator $\hat{\mathbf{T}}^{\pi}$ is defined as in (3.5) with $\mathcal{P}$ replaced by $\mathcal{P}_{n}$. The corresponding optimal empirical DR Bellman operator is $\hat{\mathbf{T}}^{*}(v):=\max_{\pi\in\Pi}\hat{\mathbf{T}}^{\pi}(v)$, defined for each $v:S\rightarrow\mathbb{R}$.
Equipped with these definitions, we present our strategy to estimate the optimal value and policy of the S-rectangular robust MDP via the empirical value function. This is motivated by the fact that $V^{*}_{\mathcal{P}}=v^{*}$ where $v^{*}$ solves (3.4).
Definition 4 (Empirical Bellman Estimators).
We define the empirical Bellman estimator $\hat{v}$ to $V^{*}_{\mathcal{P}}$ as the unique fixed point of $\hat{\mathbf{T}}^{*}$; i.e. $\hat{\mathbf{T}}^{*}(\hat{v})=\hat{v}$. Moreover, the empirical Bellman estimator for the optimal policy $\hat{\pi}^{*}$ is the maximizer of $\hat{\mathbf{T}}^{\pi}(v)$ within $\Pi$.
The remainder of the paper is devoted to theoretical analysis and numerical validation of the statistical efficiency of estimating the optimal value $V^{*}_{\mathcal{P}}=v^{*}$ using $\hat{v}$, and of estimating the optimal policy using $\hat{\pi}^{*}$. We conclude this section by introducing the following important proposition that provides an upper bound on the $l_{\infty}$ estimation error.
@@
Let $T$, $\hat{T}$ be any $\gamma$-contraction operators, and $^{*}, u^\hat{u}$ be the solution of $T(u)=u$ and $\hat{T}(u)=u$ respectively. Then, the estimation error is upper bounded by
$\|\hat{u}-u^{*}\|_{\infty}\leq\frac{1}{1-\gamma}\left\|\hat{T}(u^{*})-T(u^{*})\right\|_{\infty}$
Corollary 1.
Let $^{*},v^\hat{v}$ be the solution of $\mathcal{T}^{*}(v)=v$ and $\hat{\mathbf{T}}^{*}(v)=v$, respectively. Then, the estimation error is upper bounded by
@@
Corollary 2.
Let $V_{\mathcal{P}}^{\pi}$, $V_{\hat{\mathcal{P}}}^{\pi}$ be the solution of $v=\mathcal{T}^{\pi}(v)$ and $v=\hat{\mathbf{T}}^{\pi}(v)$ respectively. Then, the estimation error is upper bounded by
$\sup_{\pi\in\Pi}\left\|V_{\hat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\right\|_{\infty}\leq\frac{1}{1-\gamma}\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(V_{\mathcal{P}}^{\pi})-\mathcal{T}^{\pi}(V_{\mathcal{P}}^{\pi})\right\|_{\infty}$
@@ 4 Sample Complexity Bounds for the Empirical Bellman Estimator
In this section, we establish sample complexity upper bounds to achieve an absolute $\varepsilon$ error in $l_{\infty}$ distance when estimating $V^{*}_{\mathcal{P}}$ using $\hat{v}$. We focus on two specific $f$-divergence uncertainty models. When $f_{\mathrm{KL}}(t)=f(t)=t\log t$, the corresponding uncertainty set $\mathcal{P}_{s}(f_{\mathrm{KL}},\rho)$ is based on the Kullback–Leibler (KL) divergence, which is widely used in the machine learning literature. Alternatively, when $f=f_{k}$ as defined in Definition 6, the resulting $f_{k}$-divergence model captures another well-studied class of uncertainty sets [4].
@@ 4.1 The Kullback-Leibler Divergence Uncertainty Set
$\mathcal{T}^{\pi}(v)(s)=\sup_{\lambda\geq 0}\Bigg(-\lambda|\mathcal{A}|\rho-\sum_{a\in\mathcal{A}}\lambda\log\mathbb{E}_{\overline{P}_{s,a}}\left[\exp\left(-\frac{\pi(a|s)w(s,a,S)}{\lambda}\right)\right]\Bigg).$ (4.1)
where $w(s,a,S)=R(s,a,S)+\gamma v(S)$. The KL empirical DR Bellman operator $\hat{\mathbf{T}}^{\pi}$ satisfies (4.1) with $\overline{P}_{s,a}$ replaced by $\overline{P}_{s,a,n}$.
@@
$\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\|_{\infty}$ $\leq\sup_{\pi\in\Pi}\|\hat{\mathbf{T}}^{\pi}(v)-\mathcal{T}^{\pi}(v)\|_{\infty}\leq\frac{9\|R+\gamma v\|_{\infty}}{\sqrt{n\mathfrak{p}_{\wedge}}}\sqrt{\log\left(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta\right)}.$
@@
Assume the adversary chooses from a KL-divergence uncertainty set with $\rho\in(0,\infty)$. If the sample size $n\geq 12\mathfrak{p}_{\wedge}^{-1}\log(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)$, then, with probability at least $1-\eta$,
@@
Moreover, with probability at least $1-\eta$,
$\sup_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}-V_{\mathcal{P}}^{\hat{\pi}^{*}}\leq\frac{18}{(1-\gamma)^{2}\sqrt{n\mathfrak{p}_{\wedge}}}\sqrt{\log{(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)}}.$
@@
Therefore, under the KL-divergence, a total of $\widetilde{O}(|\mathcal{S}||\mathcal{A}|(1-\gamma)^{-4}\mathfrak{p}_{\wedge}^{-1}\varepsilon^{-2})$ samples from the simulator suffices to obtain an $\varepsilon$-accurate estimate of $v^{*}$ with $\hat{v}$, or an $\varepsilon$-optimal policy with high probability.
@@ 4.2 $f_{k}$-Divergence Uncertainty Set
$\mathcal{T}^{\pi}(v)(s)=-\sup_{\eta\in\mathbb{R}^{|\mathcal{A}|}}\Bigg[\sum_{a\in\mathcal{A}}\eta_{a}+c\left(\sum_{a\in\mathcal{A}}\mathbb{E}_{\overline{P}_{s,a}}\left[(\eta_{a}-\pi(a|s)w(s,a,S))_{+}^{k^{*}}\right]\right)^{1/k^{*}}\Bigg]$
where $c=c(k,\rho,|\mathcal{A}|)=|\mathcal{A}|^{1/k}\left(k(k-1)\rho+1\right)^{1/k}$, $(\cdot)_{+}=\max(\cdot,0)$ and $w(s,a,S)=R(s,a,S)+\gamma v(S)$. The $f_{k}$ empirical DR Bellman operator $\hat{\mathbf{T}}^{\pi}$ satisfies a similar equality with $\overline{P}_{s,a}$ replaced by $\overline{P}_{s,a,n}$. $\overline{P}_{s,a,n}$.
@@
Under the $f_{k}$-divergence uncertainty set with any $\rho\in(0,\infty)$, for any $v:\mathcal{S}\rightarrow\mathbb{R}$ and $n\geq 12\mathfrak{p}_{\wedge}^{-1}\log(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)$, w.p. at least $1-\eta$,
$\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\|_{\infty}\leq\sup_{\pi\in\Pi}\|\hat{\mathbf{T}}^{\pi}(v)-\mathcal{T}^{\pi}(v)\|_{\infty}\leq\frac{3\cdot 2^{k^{*}}k^{*}\|R+\gamma v\|_{\infty}}{\sqrt{n\mathfrak{p}_{\wedge}}}\sqrt{\log\left(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta\right)}.$
The proof of Proposition 3 is provided in Appendix D.1. This, combined with Proposition 1, implies the following error bound, whose proof is deferred to Appendix D.4.
@@
Assume the adversary chooses from a $f_{k}$-divergence uncertainty set with $\rho\in(0,\infty)$. If $n\geq 12\mathfrak{p}_{\wedge}^{-1}\log(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)$, then, w.p. at least $1-\eta$,
@@
Moreover, w.p. at least $1-\eta$,
$\sup_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}-V_{\mathcal{P}}^{\hat{\pi}^{*}}\leq\frac{6\cdot 2^{k^{*}}k^{*}}{(1-\gamma)^{2}\sqrt{n\mathfrak{p}_{\wedge}}}\sqrt{\log\left(4|S|^{2}|\mathcal{A}|/\eta\right)}.$
@@
Therefore, under the $f_{k}$-divergence, an $\varepsilon$-accurate estimate of $v^{*}$ with $\hat{v}$, or an $\varepsilon$-optimal policy with high probability, can be obtained using a total of $\widetilde{O}(|\mathcal{S}||\mathcal{A}|(1-\gamma)^{-4}\mathfrak{p}_{\wedge}^{-1}\varepsilon^{-2})$ samples.